OpenStudy (anonymous):

Hey Guys i really need help with this study guide its 27 questions so if osme one can help me through the whole thing or most of it id really apreciate it. oh and its Algebra 1B

OpenStudy (luckycoins888):

wat

OpenStudy (luckycoins888):

i ment what?

OpenStudy (luckycoins888):

ummmmmmmmmmmmmmmmmmm................................................

OpenStudy (luckycoins888):

what is the question?

OpenStudy (anonymous):

lemme get the picture of the triangle

OpenStudy (anonymous):

What is the side length B in the triangle below

OpenStudy (luckycoins888):

57

OpenStudy (luckycoins888):

57meters

OpenStudy (anonymous):

thats not one of the options

OpenStudy (anonymous):

625 25 22.96 527

OpenStudy (anonymous):

im really confused so if some one could actually help thatd be great

OpenStudy (radar):

You do know it has to be less than 24 right???

OpenStudy (luckycoins888):

ohhhhhh................ well................its d

OpenStudy (anonymous):

its c you have to square root it b^2=24^2-7^2

OpenStudy (radar):

pythagorean theorem rules. \[a ^{2}+b ^{2}=c ^{2}\] a = 7, c=24, now get busy and solve for c.

OpenStudy (radar):

*c was supposed to be b.

OpenStudy (texaschic101):

a^2 + b^2 = c^2 ---- c is the hypotenuse 7^2 + b^2 = 24^2 49 + b^2 = 576 b^2 = 576 - 49 b^2 = 527 b = sq rt 527 b = 22.956 rounds to 22.96

OpenStudy (anonymous):

omg i have no idea who to medal you all helped